OpenStudy (deathslasher99):

Given: d = rt, where "d" is distance, "r" is rate, and "t" is time. One car leaves point A and travels due north at 50 mi/hr. Another car leaves point A 3 hours later and travels due south at 70 mi/hr. When will the two cars be 1,000 miles apart?

OpenStudy (phi):

the easiest way to do this problem is to know that the cars are moving apart at a rate of 50+70 = 120 miles/hour

OpenStudy (deathslasher99):

ok

OpenStudy (welshfella):

yes - i used a different method and got a bit tied up!!!

OpenStudy (phi):

except for the first 3 hours, when only one car is moving at 50 mph in 3 hours it goes 150 miles 1000- 150 = 850 miles are left to go, at a rate of 120 mph

OpenStudy (deathslasher99):

im following but whats with the 120?

OpenStudy (phi):

if one car is moving north at 50 mph and the other car is moving south at 70 mph they are moving away from each other at (50+70)= 120 mph

OpenStudy (deathslasher99):

oh ok

OpenStudy (phi):

in other words, assuming they started at the same time , they would be 120 apart after 1 hour 240 miles in 2 hours or in general rate *time = 120 * # of hours.

OpenStudy (deathslasher99):

so the anwser counts as 120 mph

OpenStudy (deathslasher99):

??

OpenStudy (phi):

When will the two cars be 1,000 miles apart? you have to think it out a bit. for 3 hours , only 1 car is moving (north at 50 mph) so in 3 hours, the two cars are 150 miles apart. we want a total of 1000 miles, so they have to move another 850 miles and after 3 hours both cars are moving at a combined rate of 120 mph so we need the time it takes for them to go 850 miles then add the first 3 hours to get the total time. how long does it take to go 850 miles at rate= 120 mph ?

OpenStudy (deathslasher99):

around 7 hours, correct?

OpenStudy (phi):

7 and 1/12 hours or 7 hours 5 minutes now add the first 3 hours when only 1 car was moving the cars will be 1000 miles apart 10 hours 5 minutes after the first car started.

OpenStudy (phi):

and 7 hours 5 minutes after the 2nd car started.

OpenStudy (phi):

The question is a bit vague about how to specify the time.

OpenStudy (deathslasher99):

yes

OpenStudy (phi):

if they told us the first car left at noon, then we could say the two cars are 1000 miles apart at 10:05 pm

OpenStudy (deathslasher99):

so they where asking for the total thoe correct which is 10 hr 5m

OpenStudy (phi):

the question is not clear. I would answer the "when" by saying 10 hr 5 min after the first car started

OpenStudy (deathslasher99):

oh

OpenStudy (phi):

any questions ?

OpenStudy (deathslasher99):

nope, thank you